# Automorphisms of relatively hyperbolic groups and the Farrell–Jones Conjecture

### 23 Nov 2023 v1

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
We work with three $\phi$-invariant trees for $G$. The first is the canonical JSJ tree $T^{\mathrm{can}}$, the second is the tree $T^{\mathrm{Per}}$ obtained by applying 4.2 with ${\mathcal{H}}=\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$. The third one is obtained from $T^{\mathrm{can}}$ by blowing-up JSJ trees at QH with fibre vertices. The following lemmas motivate the construction.
@@
Since $\Psi$ and $\Theta$ fix $H$ elementwise, $\Psi$ and $\Theta$ differ by an inner automorphism in the centraliser of $H$. Since $H$ is nonelementary, its centraliser is finite (see for instance [Osi06, Theorem 4.19]). Thus, up to taking powers of $\Psi$ and $\Theta$ fixing elementwise the centraliser of $H$, we have $\Psi=\mathrm{ad}_{g}\circ\Theta$ where $g\in C_{G}(H)$ and $g\in\operatorname{\mathrm{Fix}}(\Theta)$. Thus, for every $m\geq 1$, we have $\Psi^{m}=\mathrm{ad}_{g^{m}}\circ\Theta^{m}$. As $g$ is finite order, there exists $N\in\mathbb{N}^{*}$ such that $\Psi^{N}=\Theta^{N}$. ∎
@@
    As $\Phi$ varies over the outer classes $\phi^{n}$ with $n\in\mathbb{N}^{*}$, there are only finitely many conjugacy classes of periodic subgroups $\operatorname{Per}(\Phi)$.
@@
This result does not seem surprising, and in fact the same statement is true for all hyperbolic groups (see 5.22). However, this special case is necessary to begin the arguments on JSJ decompositions we use throughout this section, including to prove the general statement.
@@
We first prove Lemma 5.2 when $G$ is a hyperbolic 2-orbifold. Recall that hyperbolic 2-orbifolds are good, and let $H$ be a characteristic finite index subgroup of $G$ corresponding to an orientable surface cover of the orbifold. (This can be obtained by taking the characteristic core of the subgroup corresponding to any such cover, since $G$ is finitely generated.) In particular, $\Phi$ preserves $H$. We now consider two periodic subgroups: $\operatorname{Per}_{G}(\Phi)\leq G$ and its subgroup $\operatorname{Per}_{H}(\Phi|_{H})\leq H$. If $g_{1}$ and $g_{2}$ are elements of $\operatorname{Per}_{G}(\Phi)$ representing the same coset of $H$ in $G$, then in fact they represent the same coset of $\operatorname{Per}_{H}(\Phi|_{H})$ in $\operatorname{Per}_{G}(\Phi)$, so this is a finite index subgroup. The restriction $\Phi|_{H}$ can be represented by an element of the mapping class group of the surface, and it follows from [Iva92] that periodic subgroups here are finitely generated. Finite generation is a commensurability invariant, so the same is true of $\operatorname{Per}_{G}(\Phi)$. Then taking a sufficiently high power to fix every element of a finite generating set shows that $\operatorname{Per}(\Phi)=\operatorname{\mathrm{Fix}}(\Phi^{k})$.
@@
Let $[g]$ be a $\phi$-periodic conjugacy class. If $g$ has finite order, then, as there exists finitely many conjugacy classes of finite order elements in $G$, some power of $\phi$ fixes $[g]$. Suppose now that $g$ has infinite order, and let $t\in\mathbb{N}^{*}$ be such that $g^{t}\in H$. Since $G$ is hyperbolic, $g^{t}$ has finitely many $t$-th roots in $G$, the number of such roots depending only on the finite numbers of orders of the finite subgroups of $G$. Thus, if $\ell\in\mathbb{N}^{*}$ is such that $\phi^{\ell}\in\operatorname{\mathrm{Out}}(H)$ fixes the conjugacy class of $g^{t}$, then a power of $\phi^{\ell}$ fixes the conjugacy class of $g$ and this power does not depend on $g$. If $H$ is a free group, then the existence (and uniformity) of $\ell$ follows from the work of Handel–Mosher [HM20, Theorem II.4.1]. If $H$ is the fundamental group of a closed orientable surface, this follows from the work of Ivanov [Iva92].
@@
Since the third statement is true for free and surface groups (by Ivanov [Iva92] for the surface case, Bestvina–Handel [BH92] for the free case with noncyclic periodic subgroups, and for instance Guirardel–Levitt [GL16] for the general case), it will suffice to bound the number of subgroups $\operatorname{Per}_{G}(\Phi)$ containing (with finite index) a given restriction $\operatorname{Per}_{H}(\Phi|_{H})$. If this is non-elementary, then it follows from Lemma 5.1 that any two automorphisms in $\phi$ fixing it have a common power, and hence the same periodic subgroups. (First replace $\Phi^{k}\operatorname{ad}(g_{1})$ and $\Phi^{\ell}\operatorname{ad}(g_{2})$ with their $\ell$-th and $k$-th powers respectively, so they represent the same outer automorphism, then another power so as to fix the common non-elementary subgroup $\operatorname{Per}_{H}(\Phi|_{H})$, then apply the lemma as written.)
@@
Now assume $\operatorname{Per}_{H}(\Phi|_{H})$ is elementary, and we want to control the periodic subgroups of $\Phi$ in $G$ restricting to it. Recall that in a hyperbolic group, every virtually cyclic subgroup is contained in a unique maximal one, and let $M$ be the maximal virtually cyclic subgroup containing $\operatorname{Per}_{H}(\Phi|_{H})$. Since $\Phi$ preserves $\operatorname{Per}_{H}(\Phi|_{H})$, it must also preserve $M$, and we consider the induced automorphism of $M$. By for instance [MO10, Lemma 6.6], $\operatorname{\mathrm{Out}}(M)$ is finite, and so passing to a power $\Phi^{k}$ the induced automorphism is inner. Composing with an inner automorphism coming from $M$, some representative $\Psi$ of $\Phi^{k}$ fixes $M$; in particular $M$ is itself a periodic subgroup.
@@
Note that this inclusion between the original periodic subgroup and the one for $\Psi$ can stay proper at all powers: the infinite order inner automorphism of $D_{\infty}$ gives an example.
@@
Let $G$ be the vertex group of a QH with finite fibre vertex. Let $F$ be a finite normal subgroup of $G_{v}$ such that $G_{v}/F$ is isomorphic to the fundamental group $\pi_{1}(\Sigma_{v})$ of a 2-orbifold $\Sigma_{v}$. Since $G_{v}$ is a hyperbolic group, it contains finitely many conjugacy classes of finite subgroups. Thus, for every $\phi\in\operatorname{\mathrm{Out}}(G_{v})$, there exists $M\geq 1$ such that $\phi^{M}$ induces an element $\phi^{M}|_{\Sigma_{v}}\in\operatorname{\mathrm{Out}}(\pi_{1}(\Sigma_{v}))$.
@@
Let $\Phi\in\operatorname{\mathrm{Aut}}(G)$. Up to taking a power of $\Phi$, we may suppose that $\Phi$ induces an element of $\operatorname{\mathrm{Aut}}(\pi_{1}(\Sigma_{v}))$. Let $\ell\in\mathbb{N}^{*}$ be the integer associated with $\Phi|_{\pi_{1}(\Sigma_{v})}$ which satisfies both Assertions $(1)$ and $(2)$. Let $g\in\operatorname{Per}(\Phi)$. Then the image of $g$ in $\pi_{1}(\Sigma_{v})$ is fixed by $\Phi^{\ell}|_{\pi_{1}(\Sigma_{v})}$. Thus, $\Phi^{\ell}$ preserves the left coset $gF$. As $F$ is finite, the automorphism $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)|}$ acts trivially on $F$. Thus, we see that $\Phi^{\ell|F||\operatorname{\mathrm{Aut}}(F)|}$ fixes $g$. This proves Assertion $(1)$. Similarly, suppose that $[g]$ is a periodic conjugacy class. Then $\Phi^{\ell}|_{\pi_{1}(\Sigma_{v})}$ preserves the conjugacy class in $\pi_{1}(\Sigma_{v})$ induced by $g$. Thus, $\Phi^{\ell}$ sends $g$ to $hk_{1}gk_{2}h^{-1}$ with $h\in G$ and $k_{1},k_{2}\in F$. Let $\Psi^{\ell}=\mathrm{ad}_{h^{-1}}\circ\Phi^{\ell}$. Then $\Psi^{\ell}$ sends $g$ to $k_{1}gk_{2}$. Note that $\Psi^{\ell|\operatorname{\mathrm{Aut}}(F)|}$ acts trivially on $F$ and sends $g$ to $k_{1}^{\prime}gk_{2}^{\prime}$ with $k_{1}^{\prime},k_{2}^{\prime}\in F$. Thus, $\Psi^{\ell|\operatorname{\mathrm{Aut}}(F)||F|}$ fixes $g$ and $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)||F|}$ fixes the conjugacy class of $g$. This proves Assertion $(2)$.
@@
For the third assertion, notice that the periodic subgroups of the induced action on $\pi_{1}(\Sigma_{v})$ contain (by passing to a finite index surface subgroup, and if necessary then an infinite index one) a preserved (and periodic) free group. This splits back to $G_{v}$, and since the arguments given earlier used only Lemma 5.1 and properties of virtually cyclic subgroups of hyperbolic groups, they apply equally well here (note that it is enough to take any non-elementary periodic subgroup to apply Lemma 5.1). Again, the final assertion follows by taking a high enough power to fix (up to composing with appropriate inner automorphisms) the generating sets of a representative of each conjugacy class. ∎
Let $v$ be a QH with fibre vertex of $T^{\mathrm{can}}$, let $e_{1},\ldots,e_{k}$ be representatives of the $G_{v}$-orbits of edges in $T^{\mathrm{can}}$ adjacent to $v$ and let $T_{v}$ be the JSJ tree of $G_{v}$ relative to $\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}([\phi^{n}|_{G_{v}}])\cup\{[G_{e_{1}}],\ldots,[G_{e_{k}}]\}$. The idea is to blow up, at every such vertex $v$ the tree $T_{v}$. But, we want the resulting tree to be compatible with $T^{\mathrm{Per}}$, so that we need to be careful when attaching the edges of $T^{\mathrm{can}}$ to vertices in $T_{v}$.
@@
Two $({\mathcal{A}},{\mathcal{P}})$-trees $T$ and $T^{\prime}$ are compatible if there exists an $({\mathcal{A}},{\mathcal{P}})$-tree $U$ such that both $T$ and $T^{\prime}$ are obtained from $U$ by collapsing some orbits of edges. By [GL17, Proposition A.26], there exists a unique such minimal tree $U$ which refines $T$ and $T^{\prime}$. The tree $U$ satisfies the following properties: a subgroup $H$ of $G$ stabilises a point in $U$ if and only if $H$ stabilises a point in both $T$ and $T^{\prime}$. Moreover, for every edge $e\in EU$, the image of $e$ in either $T$ or $T^{\prime}$ is not reduced to a point.
By universality of $T^{\mathrm{can}}$ (see [GL17, Corollary 9.18(3)]), the trees $T^{\mathrm{can}}$ and $T^{\mathrm{Per}}$ are compatible. We denote by $T^{\mathrm{ref}}_{0}$ their minimal refinement. By minimality, since $\phi$ preserves both $T^{\mathrm{can}}$ and $T^{\mathrm{Per}}$, the tree $T^{\mathrm{ref}}_{0}$ is preserved by $\phi$. The tree $T^{\mathrm{ref}}_{0}$ is obtained from $T^{\mathrm{can}}$ by blowing up, for every vertex $v\in T^{\mathrm{can}}$, a tree $S_{v}$. The tree $S_{v}$ is the minimal $G_{v}$-tree in $T^{\mathrm{ref}}_{0}$.
Suppose that $v$ is a QH vertex of $T^{\mathrm{can}}$ with finite fibre. Let $H$ be a subgroup of $G$ such that $[H]\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}([\phi^{n}|_{G_{v}}])\cup\{[G_{e_{1}}],\ldots,[G_{e_{k}}]\}$. Then $H$ is elliptic in both $T^{\mathrm{can}}$ and $T^{\mathrm{Per}}$, so that $H$ is elliptic in $T^{\mathrm{ref}}_{0}$. Since $S_{v}$ is the minimal $G_{v}$-tree in $T^{\mathrm{ref}}_{0}$, the group $H$ is also elliptic in $S_{v}$. Thus, by universality of $T_{v}$, the trees $T_{v}$ and $S_{v}$ are compatible. Let $U_{v}$ be their common minimal refinement. The tree $U_{v}$ is invariant by the outer automorphism of $G_{v}$ induced by $\phi$. Moreover, since both $S_{v}$ and $T_{v}$ are $2$-acylindrical, the minimality of $U_{v}$ implies that $U_{v}$ is $4$-acylindrical.
Let $T^{\mathrm{ref}}_{1}$ be the tree obtained from $T^{\mathrm{ref}}_{0}$ by blowing up the trees $U_{v}$ at every tree $S_{v}$ of $T^{\mathrm{ref}}_{0}$ and attaching an edge $e$ of $T^{\mathrm{ref}}_{0}$ to the centre of the fixed point set of $\operatorname{\mathrm{Stab}}(e)$ in $U_{v}$. The centre exists since for every $i\in\{1,\ldots,k\}$, the group $G_{e_{i}}$ is elliptic in $U_{v}$ and $U_{v}$ is $4$-acylindrical. It is the (necessarily common) midpoint of the longest paths in the fixed point set of $\operatorname{\mathrm{Stab}}(e)$. The centre is not necessarily a vertex of $U_{v}$, so this construction might require to subdivide some edges of $U_{v}$.
Note that the tree $T^{\mathrm{ref}}_{1}$ is preserved by $\phi$ as it is obtained from $T^{\mathrm{ref}}_{0}$ by blowing-up in a canonical way trees preserved by $\phi$.
Let $T^{\phi}_{0}$ be the tree obtained from $T^{\mathrm{ref}}_{1}$ by the following operations. We have a natural collapse map $q\colon T^{\mathrm{ref}}_{1}\to T^{\mathrm{can}}$. Let $v\in VT^{\mathrm{can}}$. If $v$ is not QH with fibre, then collapse $q^{-1}(v)$ to a point. If $v$ is QH with fibre, then $q^{-1}(v)=U_{v}$ and we collapse $U_{v}$ to the tree $T_{v}$. The resulting tree is our desired $T^{\phi}_{0}$. Note that, since $T^{\phi}_{0}$ is obtained from $T^{\mathrm{can}}$ by blowing up at each vertex $v\in VT^{\mathrm{can}}$, trees which are invariant by $\phi$ (namely, the trees $T_{v}$), the tree $T^{\phi}_{0}$ is also invariant by $\phi$.
Notice that $T^{\mathrm{ref}}_{1}$ is a common refinement of $T^{\phi}_{0}$ and $T^{\mathrm{Per}}$, so that $T^{\phi}_{0}$ and $T^{\mathrm{Per}}$ are compatible. Since $T^{\mathrm{can}}$ and all the trees $T_{v}$ are acylindrical, so is $T^{\phi}_{0}$.
Finally, let $T^{\phi}$ be the tree obtained from $T^{\phi}_{0}$ by collapsing all the edges whose endpoints are both elementary. Note that the resulting new vertices are elementary since every infinite elementary subgroup is contained in a unique maximal elementary one (see [GL15, Lemma 3.1]).
@@
The tree $T^{\phi}$ is preserved by $\phi$, it is compatible with $T^{\mathrm{Per}}$ and the action of $G$ on $T^{\phi}$ is acylindrical. Additionally, the tree $T^{\phi}$ is a bipartite tree: every edge has an endpoint which is elementary and an endpoint which is either rigid or QH with fibre. (Here, rigid vertices could correspond to rigid vertices either in $T^{\mathrm{can}}$ or in some $T_{v}$, while QH vertices are QH in a $T_{v}$, though $T_{v}$ could be a point.) Note that, unlike $T^{\mathrm{can}}$, the tree $T^{\phi}$ is not necessarily preserved by every element of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}})$. However, it is preserved by $\phi$, which is sufficient for our considerations. Moreover, the tree $T^{\phi}$ is not necessarily compatible with every $({\mathcal{A}},{\mathcal{P}})$-tree, but we will only need the fact that it is compatible with $T^{\mathrm{Per}}$. The fact that we replace $T^{\mathrm{can}}$ by $T^{\phi}$ is due to the following lemmas.

### 02 Mar 2026 v2

@@ 5.A. Trees associated with an automorphism of a one-ended relatively hyperbolic group
We work with three $\phi$-invariant trees for $G$. The first is the canonical JSJ tree $T^{\mathrm{can}}$ for the set ${\mathcal{P}}\cup{\mathcal{H}}_{0}$, the second is the tree $T^{\mathrm{Per}}$ obtained by applying ˜4.2 with ${\mathcal{H}}={\mathcal{H}}_{0}\cup\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{n})$. Note that if we restrict to $\Phi\in\operatorname{\mathrm{Aut}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0}^{(t)})$ then each $H\in{\mathcal{H}}_{0}$ is contained in some periodic subgroup for $\phi$; this explains its limited effect on the proofs below. The third one is obtained from $T^{\mathrm{can}}$ by blowing-up JSJ trees at QH with fibre vertices. The following lemmas motivate the construction.
@@
Since $\Psi$ and $\Theta$ fix $H$ elementwise, $\Psi$ and $\Theta$ differ by an inner automorphism in the centraliser of $H$. Since $H$ is nonelementary, its centraliser is finite (see for instance [45, Theorem 4.19]). Thus, up to taking powers of $\Psi$ and $\Theta$ fixing elementwise the centraliser of $H$, we have $\Psi=\mathrm{ad}_{g}\circ\Theta$ where $g\in C_{G}(H)$ and $g\in\operatorname{\mathrm{Fix}}(\Theta)$. Thus, for every $m\geq 1$, we have $\Psi^{m}=\mathrm{ad}_{g^{m}}\circ\Theta^{m}$. As $g$ is finite order, there exists $N\in\mathbb{N}^{*}$ such that $\Psi^{N}=\Theta^{N}$. ∎
@@
    As $\Phi$ varies over the outer classes $\phi^{n}$ with $n\in\mathbb{N}^{*}$, there are only finitely many conjugacy classes of maximal (for inclusion) periodic subgroups $\operatorname{Per}(\Phi)$.
@@
This result does not seem surprising, and in fact the same statement is true for all hyperbolic groups (see ˜5.24 and [28] for the torsion free case). However, this special case is necessary to begin the arguments on JSJ decompositions we use throughout this section, including to prove the general statement.
@@
We first prove Lemma 5.2 when $G$ is a hyperbolic 2-orbifold. Recall that hyperbolic 2-orbifolds are good, and let $H$ be a characteristic finite index subgroup of $G$ corresponding to an orientable surface cover of the orbifold. (This can be obtained by taking the characteristic core of the subgroup corresponding to any such cover, since $G$ is finitely generated.) In particular, $\Phi$ preserves $H$. We now consider two periodic subgroups: $\operatorname{Per}_{G}(\Phi)\leq G$ and its subgroup $\operatorname{Per}_{H}(\Phi|_{H})\leq H$. If $g_{1}$ and $g_{2}$ are elements of $\operatorname{Per}_{G}(\Phi)$ representing the same coset of $H$ in $G$, then in fact they represent the same coset of $\operatorname{Per}_{H}(\Phi|_{H})$ in $\operatorname{Per}_{G}(\Phi)$, so this is a finite index subgroup. The restriction $\Phi|_{H}$ can be represented by an element of the mapping class group of the surface, and it follows from [34] that periodic subgroups here are finitely generated. Finite generation is a commensurability invariant, so the same is true of $\operatorname{Per}_{G}(\Phi)$. Then taking a sufficiently high power to fix every element of a finite generating set shows that $\operatorname{Per}(\Phi)=\operatorname{\mathrm{Fix}}(\Phi^{k})$.
@@
Let $[g]$ be a $\phi$-periodic conjugacy class. If $g$ has finite order, then, as there exists finitely many conjugacy classes of finite order elements in $G$, some power of $\phi$ fixes $[g]$. Suppose now that $g$ has infinite order, and let $t\in\mathbb{N}^{*}$ be such that $g^{t}\in H$. Since $G$ is hyperbolic, $g^{t}$ has finitely many $t$-th roots in $G$, the number of such roots depending only on the finite numbers of orders of the finite subgroups of $G$. Thus, if $\ell\in\mathbb{N}^{*}$ is such that $\phi^{\ell}\in\operatorname{\mathrm{Out}}(H)$ fixes the conjugacy class of $g^{t}$, then a power of $\phi^{\ell}$ fixes the conjugacy class of $g$ and this power does not depend on $g$. If $H$ is a free group, then the existence (and uniformity) of $\ell$ follows from the work of Handel–Mosher [31, Theorem II.4.1]. If $H$ is the fundamental group of a closed orientable surface, this follows from the work of Ivanov [34].
@@
Since the third statement is true for free and surface groups (by Ivanov [34] for the surface case, Bestvina–Handel [12] for the free case with noncyclic periodic subgroups, and for instance Feighn–Handel [19, Corollary 11.1] for the general case), it will suffice to bound the number of subgroups $\operatorname{Per}_{G}(\Phi)$ containing (with finite index) a given restriction $\operatorname{Per}_{H}(\Phi|_{H})$. If this is non-elementary, then it follows from Lemma˜5.1 that any two automorphisms in $\phi$ fixing it have a common power, and hence the same periodic subgroups. (First replace $\Phi^{k}\operatorname{ad}(g_{1})$ and $\Phi^{\ell}\operatorname{ad}(g_{2})$ with their $\ell$-th and $k$-th powers respectively, so they represent the same outer automorphism, then another power so as to fix the common non-elementary subgroup $\operatorname{Per}_{H}(\Phi|_{H})$, then apply the lemma as written.)
@@
Now assume $\operatorname{Per}_{H}(\Phi|_{H})$ is elementary, and we want to control the periodic subgroups of $\Phi$ in $G$ restricting to it. Recall that in a hyperbolic group, every virtually cyclic subgroup is contained in a unique maximal one, and let $M$ be the maximal virtually cyclic subgroup containing $\operatorname{Per}_{H}(\Phi|_{H})$. Since $\Phi$ preserves $\operatorname{Per}_{H}(\Phi|_{H})$, it must also preserve $M$, and we consider the induced automorphism of $M$. By for instance [43, Lemma 6.6], $\operatorname{\mathrm{Out}}(M)$ is finite, and so passing to a power $\Phi^{k}$ the induced automorphism is inner. Composing with an inner automorphism coming from $M$, some representative $\Psi$ of $\Phi^{k}$ fixes $M$; in particular $M$ is itself a periodic subgroup.
@@
Note that this inclusion between the periodic subgroup for $\Phi$ and the one for $\Psi$ can stay proper at all powers. For instance take $\Phi$ to be the infinite order inner automorphism of $D_{\infty}$ and $\Psi$ to be the identity automorphism of $D_{\infty}$.
@@
Let $G_{v}$ be the vertex group of a QH with finite fibre vertex. Let $F$ be a finite normal subgroup of $G_{v}$ such that $G_{v}/F$ is isomorphic to the fundamental group $\pi_{1}(\Sigma_{v})$ of a 2-orbifold $\Sigma_{v}$. The group $F$ is the unique maximal finite normal subgroup of the hyperbolic group $G_{v}$ and is in particular characteristic. Thus, every $\phi\in\operatorname{\mathrm{Out}}(G_{v})$ induces an element $\phi|_{\Sigma_{v}}\in\operatorname{\mathrm{Out}}(\pi_{1}(\Sigma_{v}))$.
@@
Let $\Phi\in\operatorname{\mathrm{Aut}}(G_{v})$. Let $\ell\in\mathbb{N}^{*}$ be the integer associated with $\Phi|_{\pi_{1}(\Sigma_{v})}$ which satisfies both Assertions $(1)$ and $(2)$. Let $g\in\operatorname{Per}(\Phi)$. Then the image of $g$ in $\pi_{1}(\Sigma_{v})$ is fixed by $\Phi^{\ell}|_{\pi_{1}(\Sigma_{v})}$. Thus, $\Phi^{\ell}$ preserves the left coset $gF$. As $F$ is finite, the automorphism $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)|}$ acts trivially on $F$. Thus, $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)|}(g)=gf$ for some $f\in F$ fixed by $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)|}$. As $f^{|F|}=1$, we see that $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)||F|}$ fixes $g$. This proves Assertion $(1)$. Similarly, suppose that $[g]$ is a periodic conjugacy class. Then $\Phi^{\ell}|_{\pi_{1}(\Sigma_{v})}$ preserves the conjugacy class in $\pi_{1}(\Sigma_{v})$ induced by $g$. Thus, $\Phi^{\ell}$ sends $g$ to $hk_{1}gk_{2}h^{-1}$ with $h\in G_{v}$ and $k_{1},k_{2}\in F$. Let $\Psi^{\ell}=\mathrm{ad}_{h^{-1}}\circ\Phi^{\ell}$. Then $\Psi^{\ell}$ sends $g$ to $k_{1}gk_{2}$. Note that $\Psi^{\ell|\operatorname{\mathrm{Aut}}(F)|}$ acts trivially on $F$ and sends $g$ to $k_{1}^{\prime}gk_{2}^{\prime}$ with $k_{1}^{\prime},k_{2}^{\prime}\in F$. Thus, $\Psi^{\ell|\operatorname{\mathrm{Aut}}(F)||F|}$ fixes $g$ and $\Phi^{\ell|\operatorname{\mathrm{Aut}}(F)||F|}$ fixes the conjugacy class of $g$. This proves Assertion $(2)$.
@@
For the third assertion, notice that the periodic subgroups of the induced action on $\pi_{1}(\Sigma_{v})$ contain (by passing to a finite index surface subgroup, and if necessary then an infinite index one) a preserved (and periodic) free group. This splits back to $G_{v}$, and since the arguments given earlier used only Lemma˜5.1 and properties of virtually cyclic subgroups of hyperbolic groups, they apply equally well here (note that it is enough to take any non-elementary periodic subgroup to apply Lemma˜5.1). Again, the final assertion follows by taking a high enough power to fix (up to composing with appropriate inner automorphisms) the generating sets of a representative of each conjugacy class. ∎
For each non-elementary vertex $v$ of $T^{\mathrm{can}}$, let $t(v)$ be the first return power: the smallest positive power so that $\phi$ preserves the conjugacy class of $G_{v}$. Note that $G_{v}$ is self-normalising, and so $\phi^{t(v)}$ induces a well defined outer automorphism of $G_{v}$, which we denote by $\phi^{t(v)}|_{G_{v}}$. Then let
${\mathcal{H}}_{QH}=\bigcup_{\begin{subarray}{c}v\in VT^{\mathrm{can}}/G\\ v\text{ is QH}\end{subarray}}\quad\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}(\phi^{nt(v)}|_{G_{v}}).$
This set is finite by Lemma 5.2 $(3)$. Let $T^{\phi}$ be the JSJ tree of $G$ over elementary subgroups relative to ${\mathcal{P}}\cup{\mathcal{H}}_{0}\cup{\mathcal{H}}_{QH}$ given by Theorem 4.2. Note that, for every QH with fibre vertex $v\in VT^{\mathrm{can}}$ and every $H_{v}\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}([\phi^{n}|_{G_{v}}])$, there exists $H\in\bigcup_{n\in\mathbb{N}^{*}}\mathrm{Per}_{\mathrm{NP}}([\phi^{n}])$ such that $H_{v}=H\cap G_{v}$. The tree $T^{\phi}$ is $\phi$-invariant and, moreover, the tree $T^{\phi}$ is compatible with $T^{\mathrm{Per}}$.
@@
We briefly recall some facts about compatibility. Two $({\mathcal{A}},{\mathcal{P}})$-trees $T$ and $T^{\prime}$ are compatible if there exists an $({\mathcal{A}},{\mathcal{P}})$-tree $U$ such that both $T$ and $T^{\prime}$ are obtained from $U$ by collapsing some orbits of edges. By [30, Proposition A.26], there exists a unique such minimal tree $U$ which refines $T$ and $T^{\prime}$. The tree $U$ satisfies the following properties: a subgroup $H$ of $G$ stabilises a point in $U$ if and only if $H$ stabilises a point in both $T$ and $T^{\prime}$. Moreover, for every edge $e\in EU$, the image of $e$ in either $T$ or $T^{\prime}$ is not reduced to a point.
By universality of $T^{\mathrm{can}}$ (see [30, Corollary 9.18(3)]), the trees $T^{\mathrm{can}}$ and $T^{\mathrm{Per}}$ are compatible. Since the set ${\mathcal{H}}_{QH}$ consists of conjugacy classes of elliptic subgroups in both $T^{\mathrm{can}}$ and $T^{\mathrm{Per}}$, by universality of $T^{\phi}$, the trees $T^{\mathrm{can}}$, $T^{\phi}$ and $T^{\mathrm{Per}}$ are pairwise compatible.
@@
Note that, unlike $T^{\mathrm{can}}$, the tree $T^{\phi}$ is not necessarily preserved by every element of $\operatorname{\mathrm{Out}}(G,{\mathcal{P}}\cup{\mathcal{H}}_{0})$. However, it is preserved by $\phi$, which is sufficient for our considerations. Moreover, the tree $T^{\phi}$ is not necessarily compatible with every $({\mathcal{A}},{\mathcal{P}})$-tree, but we will only need the fact that it is compatible with $T^{\mathrm{Per}}$. The fact that we replace $T^{\mathrm{can}}$ by $T^{\phi}$ is due to the following lemmas.